OpenStudy (anonymous):

Solve for x in 81^x/4=1/27. can anyone help me with this?

OpenStudy (debbieg):

Can you write 1/27 as an exponential expression with a base of 3, and also write 81 the same way?

OpenStudy (debbieg):

E.g.: 81=3^4 so 81^(x/4) = \[(3^4)^{x/4}=3^x\] Now write 1/27 as 3^? same bases means that the exponents are equal. So set them equal and solve for x.